OpenStudy (anonymous):

What is a simpler form of the expression? (2n2 + 5n + 3)(4n – 5)

OpenStudy (anonymous):

\[ (2n^2 + 5n + 3)(4n - 5)\\ = 4n(2n^2 + 5n + 3) - 5(2n^2 + 5n + 3)\\ = 8n^3 + 20n^2 + 12n - 10n^2 - 25n - 15\\ = 8n^3 + 10n^2 - 13n - 15 \]

OpenStudy (anonymous):

A. 8n^3+10n^2−13n−15 I got this correct on my test.